How many gallons of pure water is to be added to 100 gallons of a 30% saline solution to make it a 10% saline solution?
Answer by stanbon(75887)

How many gallons of pure water is to be added to 100 gallons of a 30% saline solution to make it a 10% saline solution?
-------------------------------
Equation:
water + water = water
0.70*100 + x = 0.90(100+x)
----
Multiply thru by 100 to get:
70*100 + 100x = 90*100 + 90x
10x = 20*100
x = 200 gallons (amt. of pure water to be added)
===================================================
Cheers,
Stan H.
==================